OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + y = 4 x - y = 0 A.{} B.1 solution C.infinite

OpenStudy (anonymous):

well you are looking for 2 numbers, x and y when you add x and y, you want it to equal 4 when you subtract x and y, you want it to equal 0

OpenStudy (anonymous):

in essence this means that x and y must be equal to each other. simply because x - y = 0 is only true when x = y

OpenStudy (anonymous):

therefore, all you need to find is 2 numbers when added together equal 4 the only possible solution is the number 2, as 2 +2 = 4 and 2 - 2 = 0 does that make sense?

OpenStudy (anonymous):

you could solve it algebraically also x + y = 4 x - y = 0 add both equations 2x = 4 x = 2 since x = y, then y also = 2. therefore there is 1 solution (2,2)